OpenStudy (anonymous):

how do you write y=4(x-5)(x+2) in standard form

OpenStudy (anonymous):

standard form ==> Ax+By=C but since there is a x^2 y=4(x-5)(x+2) y=4(x^2+2x-5x-10) y=4(x^2-3x-10) y=?

OpenStudy (anonymous):

other references: https://answers.yahoo.com/question/index?qid=20091121134250AACbGiL

OpenStudy (anonymous):

what is y=x-5? Solution: y=x-5 y+5=x x=y+5